OpenStudy (anonymous):

There are 6 students who all want to be in a group of 3 people chosen to speak at graduation. How many different groups can be chosen?

OpenStudy (anonymous):

i'd say 10

OpenStudy (kropot72):

You need to find the number of combinations of the 6 students taken 3 at a time: \[\large 6C3=\frac{6!}{3!3!}=\frac{6\times5\times4}{3\times2\times1}=you\ can\ calculate\]

OpenStudy (anonymous):

120/6

OpenStudy (anonymous):

?

OpenStudy (kropot72):

Yes, 20 is correct.

OpenStudy (anonymous):

really? i did it by writing it. the first time going across: 1,2,3,4,5,6= 4 ways. did it again and got 3 ways. again and got 2 times. last i got 1. then i added them to get 10

OpenStudy (anonymous):

so final answer is 20?

OpenStudy (anonymous):

123 and 456,124 and 356, 125 and 346. then 234 and 156, 235 and 146, 236 and 145. then 345 and 126. so i got 7? confused

OpenStudy (kropot72):

The number of combinations of n different things taken r at a time can be written nCr.\[\large nCr=\frac{n(n-1)(n-2)\ .........\ (n-r+1)}{1\times2\times3\ ...... \times r}=\frac{n!}{r!(n-r)!}\]

OpenStudy (kropot72):

Yes, the number of combinations of 6 students taken 3 at a time is 20. Therefore 20 different groups can be chosen.

OpenStudy (anonymous):

alright. i'd go with what he said

OpenStudy (anonymous):

can you help me with this one?

OpenStudy (anonymous):

you have 20 books and want to make a reading list that has 3 books on it. How many different reading lists can you make? (2 reading lists are different if they have the same books but in a different order

OpenStudy (anonymous):

6,840 i think if i follow crept correctly

OpenStudy (anonymous):

kropot. think my computer "corrected" it to crept

OpenStudy (kropot72):

np! In the question with the books, the number of permutations of 20 different books taken 3 at a time is required. \[\large 20P3=\frac{20!}{(20-3)!}=20\times19\times18\]

OpenStudy (anonymous):

6,840? correct

OpenStudy (anonymous):

kropot is correct with the equation which would equal to 6840